OpenStudy (trwatkins1):

please help Points V, W, X, Y, and Z are graphed on a number line. The coordinate of point V is –8. What points are at a distance of 5 units from point V? Choose all answers that are correct. A. W = –13 B. X = –8 C. Y = –3 D. Z = 3

OpenStudy (anonymous):

is there any other information or a picture given?

OpenStudy (trwatkins1):

nope

OpenStudy (trwatkins1):

@Spring98

OpenStudy (anonymous):

ok then based on that i would say a because im assuming they fall in alphabetical order and -13 is 5 away from -8

OpenStudy (trwatkins1):

so you think a?

OpenStudy (anonymous):

correct

OpenStudy (trwatkins1):

@jeweaver145 you do know its multiple choice right?

OpenStudy (anonymous):

yes i do, which is why i said A was correct.

OpenStudy (trwatkins1):

is that the only one?

OpenStudy (trwatkins1):

?

OpenStudy (anonymous):

If V = -8, add and subtract 5 to get the points 5 away from it. -8 + 5 = -8 - 5 =

OpenStudy (trwatkins1):

so a & c?

OpenStudy (trwatkins1):

@peachpi

OpenStudy (anonymous):

yes